IR SPECTROSCOPY OF LONG CARBON CHAINS AND THEIR DERIVATIVES IN CRYOGENIC MATRI-CES: SPECIES IDENTIFICATION BY ^{13}C_{3}→^{12}C_{3} ISOTOPIC SUBSTITUTION.

\underline{DMITRY STRELNIKOV} and WOLFGANG KRÄTSCHMER, Max-Planck-Institut für Kernphysik, D-69029 Heidelberg, PO Box 103980, Germany.

Along with matrix gas, we co-deposited carbon vapor from two separated carbon evaporation sources, namely a ^{12}Csource and a ^{13}C-source. Under the applied conditions the C_{3} molecules are the most abundant species in both carbon sources. We observed that in the formation process of long carbon chains the C_{3} units play a major role. This finding was used to establish a new species identification method based on isotopic replacement. The old technique of atomic 13C→^{12}C substitution can be reasonably applied only for molecules with a small number of carbon atoms (n9). Using the new method of molecular ^{13}C_{3}→^{12}C_{3} substitution, quite simple distributions having small number of isotopomeric IR absorptions were obtained also for larger species. We present data for C_{n} and C_{n}O_{1,2} (n6).